# Measurement of the longitudinal flow-plane decorrelation using multi-plane cumulants in $\sqrt{s_{{}_{\mathrm{NN}}}}$ = 200 GeV Au+Au, Ru+Ru, and Zr+Zr collisions

### v1

@@ Abstract
Measurements of the variation of anisotropic flow-plane angles ($\Psi_{n}$) with rapidity, commonly known as the flow-plane decorrelation, provide important insights into the initial conditions of the matter produced in heavy-ion collisions. In this paper, using data collected by the STAR experiment, we report the first measurement of the four-plane cumulant observable $T_{n}\{ba;dc\}=\langle\langle\sin[n(\Psi^{b}_{n}-\Psi^{a}_{n})]\sin[n(\Psi^{d}_{n}-\Psi^{c}_{n})]\rangle\rangle$, where superscripts $a$, $b$, $c$, and $d$ denote sequential regions in pseudorapidity with $a$ corresponding to the most backward pseudorapidity region. The measurement is performed for the elliptic and triangular flow (i.e. $n=2$ and $3$) in Au+Au and isobar (Ru+Ru, Zr+Zr) collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}$ = 200 GeV. The goal of calculating the correlation of the flow-plane angle variations, from backward pseudorapidity ($\eta_{a}$) to mid-pseudorapidity ($\eta_{b}$) region and from mid-pseudorapidity $(\eta_{c}$) to forward pseudorapidity ($\eta_{d}$) region, is to probe the systematic variation of flow angle over a wide pseudorapidity range. In mid-central collisions ($10-30\%$ centrality), we find $T_{2}\{ba;dc\}=-0.004\pm 0.001(\rm stat)\pm 0.002(\rm syst)$ independent of the collision system. Such a small value of $T_{2}$ favors a “random-walk” variation of the flow-plane angles, where the rapidity correlation length is smaller than the pseudorapidity region under study. These measurements provide new information on the decorrelation patterns in the system and offer a quantitative estimate of possible systematic variations in anisotropic flow angles-such as “twist” between forward and backward pseudorapidity regions. This opens new opportunities for understanding the three-dimensional structure of the quark-gluon plasma created in heavy-ion collisions.
@@ I Introduction
The STAR experiment at the Relativistic Heavy Ion Collider (RHIC) at Brookhaven National Laboratory (BNL) focuses on investigating the formation of a deconfined state of quarks and gluons known as the quark-gluon plasma (QGP) et al. (STAR Collaboration); Busza et al. (2018). One of the primary goals of STAR and other heavy ion experiments is to probe the mechanism behind the QGP formation and to understand its evolution and hadronization processes. It has long been realized that the QGP created in such collisions exhibits significant initial-state spatial anisotropies which originate from both the geometry of non-central nuclear collisions and the probabilistic nature of nucleon-nucleon interactions. The spatial anisotropies in the initial-state matter distribution lead to momentum anisotropies of the final-state particles due to the space-momentum interactions during system evolution.
@@
The reaction plane angle $\Psi_{\rm RP}$ is defined by the direction of the impact parameter vector. Using $\Psi_{\rm RP}$ in place of $\Psi_{n}$ in the decomposition Eq. 1 for all harmonics, corresponds to the flow ($v_{n}$) definition in the reaction-plane anisotropic flow picture. While this picture is well defined theoretically, in practice it is not very useful, as the reaction-plane angle can not be measured experimentally. In most cases, the flow itself (e.g., anisotropic particle distribution) is used to determine the flow-symmetry planes $\Psi_{n}$ as those corresponding to the maximum amplitude of a given harmonic. Due to the initial-state geometry fluctuations, flow planes $\Psi_{n}$ in general do not coincide with the reaction plane Alver and Roland (2010). In this so-called participant picture of flow, the flow development is determined by the initial geometry of the system, including fluctuations caused by the probabilistic character of the interactions of nucleons inside the colliding nuclei. The participant-plane description is meaningful only if the flow angles depend relatively weakly on the particle rapidity and, more generally, on the particle transverse momentum ($p_{\mathrm{T}}$) as well as particle type. Otherwise, the Fourier decomposition in Eq. 1 becomes impractical.
@@
The study of the flow-plane angle $\Psi_{n}$, along with its fluctuations and correlations across rapidity, is of particular interest. Equally important are similar measurements of the flow coefficients $v_{n}$. These quantities are determined by fluctuations in the initial matter distribution. Consequently, they provide insights into various aspects of nuclear interactions, such as baryon stopping, shadowing, and sub-nucleon structure. Such fluctuations or variations of the flow plane with rapidity are referred to as longitudinal flow-plane decorrelations.
@@
For discussion purposes, it is useful to separate the different sources of flow-plane decorrelations in rapidity. Short-range effects, such as jet energy deposition, influence only a limited region in rapidity. Long-range effects, on the other hand, arise from correlations like those in the “twisted fireball” scenario Bożek et al. (2011); Behera et al. (2020), where the flow plane in the forward region is primarily determined by the participant distribution from the projectile nucleus, while at backward rapidity, it is governed by the participants of target nucleus. The short-range effects when measured on a distance scale much larger than the correlation length appear as a “random-walk” variations in the flow angle. The resulting decorrelation is referred to as a random tilt. The “twisted fireball” scenario, on the other hand, can be visualized as a monotonic change in the flow-plane angle with rapidity. Another scenario corresponding to the long-range variation of flow plane occurs when the flow-plane angles in the forward and backward regions are close to each other, while those at the central region experiences relatively large fluctuations. This case would correspond to a non-monotonic change in the flow-plane angle with rapidity. See Refs. Pang et al. (2015, 2016); Wu et al. (2019); Jia and Huo (2014a, b); Xiao et al. (2013); Jia et al. (2024) for a more detailed discussion on the mechanisms of the flow-plane decorrelations.
@@
$\begin{split}r_{n}(\eta_{a\rightarrow c/b})&\equiv\frac{\langle\cos[n(\phi^{c}-\Psi_{n}^{a})]\rangle}{\langle\cos[n(\phi^{b}-\Psi_{n}^{a})]\rangle}\\ &=\frac{\langle v_{n}(\eta_{c})\cos[n(\Psi_{n}^{c}-\Psi_{n}^{a})]\rangle}{\langle v_{n}(\eta_{b})\cos[n(\Psi_{n}^{b}-\Psi_{n}^{a})]\rangle}\\ &\approx\frac{\langle\cos[n(\Psi_{n}^{c}-\Psi_{n}^{a})]\rangle}{\langle\cos[n(\Psi_{n}^{b}-\Psi_{n}^{a})]\rangle}\\ &\equiv r^{\Psi}_{n}(\eta_{a\rightarrow c/b}),\end{split}$ (2)
@@
where delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the average over events. $\phi^{b}$ and $\phi^{c}$ are the azimuthal angles of particles from the two symmetric mid-pseudorapidity windows $b$ and $c$, respectively, on the backward and forward regions (see Fig. 1). Similarly, $\Psi_{n}^{a}$, $\Psi_{n}^{b}$, and $\Psi_{n}^{c}$ are the flow-plane angles of $n^{\rm th}$-order harmonic flow obtained from the backward pseudorapidity reference window $a$ ($\eta_{a}<0$) and two mid-pseudorapidity windows $b$ ($\eta_{b}<0$) and $c$ ($\eta_{c}>0$), respectively. The notations $\eta_{a}$, $\eta_{b}$, and $\eta_{c}$ refer to the ranges in pseudorapidity of three windows such that the reference window $a$ in the backward region is sufficiently far from the two mid-pseudorapidity windows $b$ and $c$ to introduce large pseudorapidity gaps, $|\eta_{b}-\eta_{a}|$ and $|\eta_{c}-\eta_{a}|$, in the measurement. The reference window in Eq. 2 can also be selected from the forward region (such as window $d$ in Fig. 1), provided that the mid-pseudorapidity regions $b$ and $c$ are simultaneously interchanged in the definition.
@@
In symmetric collisions, such as Au+Au or Ru+Ru (Zr+Zr), the flow coefficients $v_{n}(\eta_{b})$ and $v_{n}(\eta_{c})$, corresponding to symmetric negative and positive pseudorapidity ranges $\eta_{b}$ and $\eta_{c}$, respectively, are expected to be identical when averaged over all events. Hence, the deviation of the ratio in Eq. 2 from unity would indicate the presence of flow decorrelation in the collisions. The observable $r_{n}(\eta_{a\rightarrow c/b})$ in Eq. 2 includes contributions from the decorrelations in the flow angles ($r_{n}^{\Psi}(\eta_{a\rightarrow c/b})$) as well as in the flow magnitudes, which is the reason for an approximate sign in the third line. As shown in Ref. Bozek and Broniowski (2018), the decorrelations in the flow magnitudes are relatively small, at the level of a few percent. However, the effect of using the flow magnitude weight in the calculation of the flow-plane decorrelation can be rather significant and comparable to the flow-angle decorrelation itself. This is due to the fact that the flow-plane fluctuations are smaller in the events with larger flow magnitude.
@@
The observable $r_{n}(\eta_{a\rightarrow c/b})$ is sensitive to both the random tilt and the long-range systematic rotation of flow planes as in the “twisted fireball” scenario. As shown in Ref. Xu et al. (2022), the ratio $r_{n}^{\Psi}(\eta_{a\rightarrow c/b})$ is independent of the reference rapidity window $a$ in the random-walk scenario of the flow-plane angle variations.
In the small angle approximation, the observable $r_{n}^{\Psi}(\eta_{a\rightarrow c/b})$ can be expressed as
@@
$\begin{split}&r^{\Psi}_{n}(\eta_{a\rightarrow c/b})\\ &\approx 1-\frac{n^{2}}{2}\langle(\Delta\Psi_{n}^{b\rightarrow c})^{2}\rangle-n^{2}\langle\Delta\Psi_{n}^{a\rightarrow b}\cdot\Delta\Psi_{n}^{b\rightarrow c}\rangle,\end{split}$ (3)
@@
The observable defined in Eq. 2 is designed to measure the relative decrease in the anisotropic flow caused by the $\Psi_{n}$ decorrelation. This is evaluated over a “$2\eta$” distance (where $|\eta_{b}|=|\eta_{c}|=\eta$) between regions $b$ and $c$, with respect to a reference pseudorapidity window $a$. In practice, the decorrelation measured by Eq. 2 is independent of the reference pseudorapidity window $a$ only if there is no contribution from the term $\langle\Delta\Psi_{n}^{a\rightarrow b}\cdot\Delta\Psi_{n}^{b\rightarrow c}\rangle$ (see Eq. 3), i.e., if the consecutive changes in the flow-plane orientations $\Delta\Psi_{n}^{a\rightarrow b}$ and $\Delta\Psi_{n}^{b\rightarrow c}$ of the flow-plane angles are independent.
This paper presents an investigation of the longitudinal flow-plane decorrelations using a four-plane cumulant $T_{n}$ observable, proposed in Ref. Xu et al. (2022), using flow-plane angles from the backward, forward, and two mid-pseudorapidity regions. The $T_{n}$ observable is defined by
$T_{n}\{ba;dc\}=\langle\langle\,\sin[n(\Psi_{n}^{b}-\Psi_{n}^{a})]\sin[n(\Psi_{n}^{d}-\Psi_{n}^{c})]\,\rangle\rangle,$ (5)
[…]
The $T_{n}$ observable is largely insensitive to correlations unrelated to the flow-plane orientation (i.e., non-flow correlations due to few-particle clusters such as jets, heavy resonance decays, local charge conservation, etc.), as it is based on the fourth-order cumulant Borghini et al. (2001); Voloshin et al. (2010); et al. (2002). For the same reason, it is also insensitive to the “random-walk” type of flow decorrelations. Global momentum conservation, although it does not contribute to harmonic correlations with $n>1$, arises from many different few-particle processes and is therefore also suppressed in higher-order cumulant calculations. Similarly, the ratio observable $r_{n}(\eta_{a\rightarrow c/b})$ in Eq. 2 is sensitive to the flow magnitude fluctuations. As the flow-plane fluctuations are smaller in events with the larger flow Bozek and Broniowski (2018), the measurements of $r_{n}(\eta_{a\rightarrow c/b})$ and $T_{n}$ are likely to be biased toward smaller values.
@@
where $\Delta\Psi^{a\rightarrow b}_{n}=\Psi_{n}^{b}-\Psi_{n}^{a}$ and $\Delta\Psi^{c\rightarrow d}_{n}=\Psi_{n}^{d}-\Psi_{n}^{c}$ are $\Psi_{n}$ twists from the pseudorapidity subevents $a\rightarrow b$ and $c\rightarrow d$. Therefore, $T_{n}$ represents the covariance between changes in orientation of flow plane from pseudorapidity window $a$ to $b$ and that from $c$ to $d$, quantifying the overall azimuthal twist of flow plane on moving forward in the pseudorapidity. One would expect a positive value of $T_{n}\{ba;dc\}$ when the flow-plane angles at backward and forward pseudorapidities are fluctuating in opposite directions to that at mid-pseudorapidity. On the other hand, the negative sign of $T_{n}\{ba;dc\}$ would correspond to the decorrelation pattern when the flow-plane angles at backward and forward pseudorapidities fall on the same direction relative to that at mid-pseudorapidity. These two distinct cases $T_{n}>0$ and $T_{n}<0$ are identified as the “S-shaped” and “C-shaped” decorrelations, respectively, in Ref. Xu et al. (2022).
In this analysis, we attempt to determine and identify the dominant pattern of the longitudinal flow-plane decorrelations by measuring $r_{n}$, $r_{n}^{\Psi}$, and $T_{n}$ observables using the flow-plane angles reconstructed in the forward, backward, and mid-pseudorapidity regions in isobar (Ru+Ru, Zr+Zr) and Au+Au collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV with the STAR detector at RHIC.

### v2

@@ Abstract
Measurements of the variation of anisotropic flow-plane angles ($\Psi_{n}$) with rapidity, commonly known as the flow-plane decorrelation, provide important insights into the initial conditions of the matter produced in heavy-ion collisions. In this paper, using data collected by the STAR experiment, we report the first measurement of the four-plane correlator observable $T_{n}\{ba;dc\}=\langle\langle\sin[n(\Psi^{b}_{n}-\Psi^{a}_{n})]\sin[n(\Psi^{d}_{n}-\Psi^{c}_{n})]\rangle\rangle$, where superscripts $a$, $b$, $c$, and $d$ denote sequential pseudorapidity ($\eta$) regions with $a$ corresponding to the most backward region, $b$ and $c$ close to mid-rapidity with $\eta_{b}<0$ and $\eta_{c}>0$, and $d$ being the most forward. The measurement is performed for the elliptic and triangular flow (i.e. $n=2$ and $3$) in Au+Au and isobar (Ru+Ru, Zr+Zr) collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}$ = 200 GeV. The goal of calculating the correlation of the flow-plane angle variations from backward to mid-central, and from mid-central to forward regions, is to probe the systematic variation of flow angle over a wide $\eta$ range. In mid-central collisions ($10-30\%$ centrality), we find $T_{2}\{ba;dc\}=-0.004\pm 0.001(\rm stat)\pm 0.002(\rm syst)$ independent of the collision system. Such a small value of $T_{2}$ favors a “random-walk” variation of the flow-plane angles, where the rapidity correlation length is smaller than the entire region under study. These measurements provide new information on the decorrelation patterns in the system and offer a quantitative estimate of possible systematic variations in anisotropic flow angles such as “twist” between forward and backward regions. This opens new opportunities for understanding the three-dimensional structure and the time evolution of the quark-gluon plasma created in heavy-ion collisions.
@@ I Introduction
The STAR experiment at the Relativistic Heavy Ion Collider (RHIC) at Brookhaven National Laboratory (BNL) focuses on investigating the formation of a deconfined state of quarks and gluons known as the quark-gluon plasma (QGP) Collaboration) (2005); Busza et al. (2018). One of the primary goals of STAR and other heavy ion experiments is to probe the mechanism behind the QGP formation and to understand its evolution and hadronization processes. It has long been realized that the QGP created in such collisions exhibits significant initial-state spatial anisotropies which originate from both the geometry of non-central nuclear collisions and the probabilistic nature of nucleon-nucleon interactions. The spatial anisotropies in the initial-state matter distribution lead to momentum anisotropies of the final-state particles due to the space-momentum correlations during system evolution.
@@
The reaction plane angle $\Psi_{\rm RP}$ is defined as a plane spanned by the impact parameter vector and the beam axis. Using $\Psi_{\rm RP}$ in place of $\Psi_{n}$ in the decomposition Eq. 1 for all harmonics, corresponds to the flow ($v_{n}$) definition in the reaction-plane anisotropic flow picture. While this picture is well defined theoretically, in practice it is not very useful, as the reaction-plane angle can not be measured experimentally. In most cases, the flow itself (e.g., anisotropic particle distribution) is used to determine the flow-symmetry planes $\Psi_{n}$ as those corresponding to the maximum amplitude of a given harmonic Voloshin et al. (2010). Due to the initial-state geometry fluctuations, flow planes $\Psi_{n}$ in general do not coincide with the reaction plane Alver and Roland (2010). In this so-called participant picture of flow, the flow development is primarily driven by the initial geometry of the system, including fluctuations caused by the probabilistic character of the interactions of nucleons inside the colliding nuclei. The participant-plane description is meaningful only if the flow angles depend relatively weakly on the particle rapidity and, more generally, on the particle transverse momentum ($p_{\mathrm{T}}$) as well as particle type. Otherwise, the Fourier decomposition in Eq. 1 is not practically useful.
@@
The study of the flow-plane angle $\Psi_{n}$ variation in longitudinal direction, along with its fluctuations and correlations in $\eta$, is of particular interest. Together with similar measurements of the flow coefficients $v_{n}$ they provide insights into various aspects of nuclear interactions, such as baryon stopping, shadowing, and sub-nucleon structure.
@@
For discussion purposes, it is useful to separate the different sources of flow-plane decorrelations. Short-range effects, such as jet energy deposition, influence only a limited region in pseudorapidity. Long-range effects, on the other hand, arise from correlations like those in the “twisted fireball” scenario Bożek et al. (2011); Behera et al. (2020), where the flow plane in the forward region is primarily determined by the participant distribution from the projectile nucleus, while at backward pseudorapidity, it is governed by the participants of target nucleus. The short-range effects when measured on a distance scale much larger than the correlation length appear as a “random-walk” variations in the flow angle. The resulting decorrelation is referred to as a random tilt. The “twisted fireball” scenario, on the other hand, can be visualized as a monotonic change in the flow-plane angle with pseudorapidity. Another scenario corresponding to the long-range variation of flow plane occurs when the flow-plane angles in the forward and backward regions are close to each other, while those at the central region experiences relatively large fluctuations. This case would correspond to a non-monotonic change in the flow-plane angle with pseudorapidity. See Refs. Pang et al. (2015, 2016); Wu et al. (2019); Jia and Huo (2014b, a); Xiao et al. (2013); Jia et al. (2024) for a more detailed discussion on the mechanisms of the flow-plane decorrelations.
@@
$\begin{split}r_{n}(\eta_{a\rightarrow c/b})&\equiv\frac{\langle\cos[n(\phi^{c}-\Psi_{n}^{a})]\rangle}{\langle\cos[n(\phi^{b}-\Psi_{n}^{a})]\rangle}\\ &=\frac{\langle v_{n}(\eta_{c})\cos[n(\Psi_{n}^{c}-\Psi_{n}^{a})]\rangle}{\langle v_{n}(\eta_{b})\cos[n(\Psi_{n}^{b}-\Psi_{n}^{a})]\rangle}\end{split}$ (2)
@@
where $\langle$$\rangle$ denotes the average over events. $\phi^{b}$ and $\phi^{c}$ are the azimuthal angles of particles from the two symmetric mid-pseudorapidity windows $b$ and $c$, respectively, on the backward and forward regions (see Fig. 1). Similarly, $\Psi_{n}^{a}$, $\Psi_{n}^{b}$, and $\Psi_{n}^{c}$ are the flow-plane angles of $n^{\rm th}$-order harmonic flow obtained from the reference $\eta$ window $a$ ($\eta_{a}<0$), and two symmetric windows $b$ ($\eta_{b}<0$) and $c$ ($\eta_{c}>0$). The notations $\eta_{a}$, $\eta_{b}$, and $\eta_{c}$ refer to the ranges in pseudorapidity of three windows such that the reference window $a$ is sufficiently far from windows $b$ and $c$ to introduce large pseudorapidity gaps, $|\eta_{b}-\eta_{a}|$ and $|\eta_{c}-\eta_{a}|$, in the measurement, ensuring suppression of non-flow. The reference window in Eq. 2 can also be selected from the forward region (such as window $d$ in Fig. 1), provided that regions $b$ and $c$ are simultaneously interchanged in the definition.
@@
In symmetric collisions, such as Au+Au or Ru+Ru (Zr+Zr), the flow coefficients $v_{n}(\eta_{b})$ and $v_{n}(\eta_{c})$, corresponding to symmetric negative and positive ranges $\eta_{b}$ and $\eta_{c}$, respectively, are expected to be identical when averaged over all events. Hence, the deviation of the ratio in Eq. 2 from unity would indicate the presence of flow decorrelation in the collisions. The observable $r_{n}(\eta_{a\rightarrow c/b})$ in Eq. 2 includes contributions from the decorrelations in the flow angles as well as in the flow magnitudes. As shown in Ref. Bozek and Broniowski (2018), the decorrelations in the flow magnitudes are relatively small, at the level of a few percent. However, the effect of using the flow magnitude weight in the calculation of the flow-plane decorrelation can be rather significant and comparable to the flow-angle decorrelation itself. This is due to the fact that the flow-plane fluctuations are smaller in the events with larger flow magnitude Bozek and Broniowski (2018). Experimentally, the flow (event) planes are determined from flow itself. This determination is more accurate in events with larger flow, and thus the measurements of flow-plane decorrelations always implicitly include flow magnitude weights. Below, in the discussion of the results of our measurements in terms of flow-plane decorrelations we assume that the corresponding average values are obtained with flow magnitude weights.
Figure 1: Diagram illustrating four pseudorapidity regions used for $T_{n}$ calculations.
@@
The observable $r_{n}(\eta_{a\rightarrow c/b})$ is sensitive to both the random tilt and the long-range systematic rotation of flow planes as in the “twisted fireball” scenario. As shown in Ref. Xu et al. (2022), the ratio $r_{n}(\eta_{a\rightarrow c/b})$ is independent of the reference pseudorapidity window $a$ in the random-walk scenario of the flow-plane decorrelations.
In the small angle approximation, the observable $r_{n}(\eta_{a\rightarrow c/b})$ can be expressed as
@@
$\begin{split}&r_{n}(\eta_{a\rightarrow c/b})\\ &\approx 1-\frac{n^{2}}{2}\langle(\Delta\Psi_{n}^{b\rightarrow c})^{2}\rangle-n^{2}\langle\Delta\Psi_{n}^{a\rightarrow b}\cdot\Delta\Psi_{n}^{b\rightarrow c}\rangle,\end{split}$ (3)
@@
The observable defined in Eq. 2 is designed to measure the relative decrease in the anisotropic flow caused by the $\Psi_{n}$ decorrelation over a $\eta_{c}-\eta_{b}$ distance. In practice, the decorrelation measured by Eq. 2 is independent of the reference window $a$ only if there is no contribution from the term $\langle\Delta\Psi_{n}^{a\rightarrow b}\cdot\Delta\Psi_{n}^{b\rightarrow c}\rangle$ (see Eq. 3), i.e., if the consecutive changes in the flow-plane orientations $\Delta\Psi_{n}^{a\rightarrow b}$ and $\Delta\Psi_{n}^{b\rightarrow c}$ of the flow-plane angles are independent.
This paper presents a study of the longitudinal flow-plane decorrelations using a four-plane $T_{n}$ correlator, proposed in Ref. Xu et al. (2022), using the flow-plane angles from the backward, forward, and two mid-rapidity regions. The $T_{n}$ observable is defined by
$T_{n}\{ba;dc\}$ $\equiv$ $\langle\langle\,\sin[n(\Psi_{n}^{b}-\Psi_{n}^{a})]\sin[n(\Psi_{n}^{d}-\Psi_{n}^{c})]\,\rangle\rangle$ (5)
$=$ $\langle\sin[n(\Psi_{n}^{b}-\Psi_{n}^{a})]\sin[n(\Psi_{n}^{d}-\Psi_{n}^{c})]\rangle$
$-\frac{1}{2}\langle\cos[n(\Psi_{n}^{b}-\Psi_{n}^{d})]\rangle\langle\cos[n(\Psi_{n}^{a}-\Psi_{n}^{c})]\rangle$
$+\frac{1}{2}\langle\cos[n(\Psi_{n}^{a}-\Psi_{n}^{d})]\rangle\langle\cos[n(\Psi_{n}^{b}-\Psi_{n}^{c})]\rangle,$
where double angle brackets $\langle\langle$$\rangle\rangle$ denote the cumulant part, obtained by subtracting from the 4-plane correlator (first line in Eq. I) the contributions from all lower-order correlations (subsequent lines). $\Psi^{i}_{n}\textrm{ (for }i=a,b,c,d)$ are the flow-plane angles of $n^{\rm th}$-order harmonic flow. The superscript in the flow-plane angle denotes subevents, $a$ through $d$, as illustrated in Fig. 1. In the expansion Eq. I we do not show the terms that are present in the full cumulant expansion of the four-plane correlator but that are equal to zero due to azimuthal symmetry (e.g. proportional to averages of sine or cosine of a single flow-plane angle; see Ref. Xu et al. (2022) for details). In the analysis of the real data, to account for small anisotropies in particle detection the vanishing of such terms is achieved by recentering of the flow vectors, see below.
The main property of the observable $T_{n}$ is that it is nonzero only if all four flow-plane angles are correlated. Thus it is largely insensitive to correlations unrelated to the flow-plane orientation (i.e., non-flow correlations due to few-particle clusters such as jets, heavy resonance decays, local charge conservation, etc.), which are suppressed similarly to that in flow measurements based on fourth-order cumulants Borghini et al. (2001); Voloshin et al. (2010); et al. (2002). The AMPT studies of $T_{n}$ observable Xu et al. (2022) confirm this conclusion. For the same reason, $T_{n}$ is also insensitive to the “random-walk” type of flow decorrelations. Similarly to the ratio observable $r_{n}(\eta_{a\rightarrow c/b})$, the exprimental measurements of $T_{n}$ are sensitive to the flow magnitude fluctuations and effectively include flow magnitudes as a weight.
@@
where $\Delta\Psi^{a\rightarrow b}_{n}=\Psi_{n}^{b}-\Psi_{n}^{a}$ and $\Delta\Psi^{c\rightarrow d}_{n}=\Psi_{n}^{d}-\Psi_{n}^{c}$. Therefore, $T_{n}$ represents the covariance between changes in orientation of flow plane from window $a$ to $b$ and that from $c$ to $d$, quantifying the overall azimuthal twist of flow plane on moving forward in the pseudorapidity. One would expect a positive value of $T_{n}\{ba;dc\}$ when the flow-plane angles at backward and forward regions are fluctuating in opposite directions to that at mid-rapidity. On the other hand, the negative sign of $T_{n}\{ba;dc\}$ would correspond to the decorrelation pattern when the flow-plane angles at backward and forward regions fall on the same direction relative to that at the central region. These two distinct cases $T_{n}>0$ and $T_{n}<0$ are identified as the “S-shaped” and “C-shaped” decorrelations, respectively, in Ref. Xu et al. (2022).
In this analysis, we attempt to determine and identify the dominant pattern of the longitudinal flow-plane decorrelations by measuring $r_{n}$ and $T_{n}$ observables using the flow-plane angles reconstructed in the forward, backward, and mid-pseudorapidity regions in isobar (Ru+Ru, Zr+Zr) and Au+Au collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV with the STAR detector at RHIC.